# Towards Pervasive Distributed Agentic Generative AI - A State of The Art

### v1

@@ 4.3.3. Action
  • Tool Use / API Calls: Agents can leverage external services or system functionalities by using specific tools or making API calls. This includes a wide range of interactions, such as operating smartphone applications, sending emails, querying databases, retrieving weather information, or controlling external devices. For example, in (Wen et al., 2024a), the agents possesses a defined set of possible actions (CLICK, CHECK/UNCHECK, SCROLL¡DIRECTION¿, INPUT¡TEXT¿) to interact with a smartphone device’s interface. In (Rivkin et al., 2024), the agent flexibly manages smart home devices using a dedicated device interaction tool. The API calls are executed using the SmartThings REST API for reading device attributes and sending commands, with an automatic error correction mechanism in case of call failures. Similarly, in (Li et al., 2024b) the agent is equipped with a series of tools, including Knowledge Retrieval, Data Extraction, Classification, and Long-Term Memory Retrieval, which it can use for intrusion detection analysis. These tools are invoked using a JSON format that specifies an action name (the tool’s identifier) and an action input (the associated settings or parameters for that tool).
@@
  • Code Execution: Agents may generate and execute code snippets to test conditions or directly interact with underlying systems. For example, in (Wen et al., 2024b), the agent interacts with the environment by translating a user’s natural language task into executable Python-like scripts. These scripts are generated by a small language model based on a structured app document and are then executed on the device to manipulate page contents elements, such as tapping buttons or scrolling content. Also, in (Rao et al., 2024), a large language model is employed to generate Python code that guides the placement of microservices between edge and cloud infrastructure, based on real-time workload and latency data. This enables the LLM to adaptively recommend whether a microservice should run on the edge or in the cloud, with new placement decisions being automatically integrated into the running application.
@@
  • User Messaging (Natural Language Interaction): Agents can also act by directly interacting with users through natural language messages, providing support as intelligent assistants. For instance, in (Fan et al., 2025), the agent communicates its analysis of manufacturing images to users, performing anomaly detection and scene understanding to provide comprehensive support to smart factory employees. Also, in (Zagar et al., 2025), the system uses a unified interface to interact with users naturally via chat, enabling secure, real-time processing of health data in healthcare domains such as Electronic Health Record (EHR) analysis, patient data explanation, and medical form automation.
4.4. Applications and Evaluations of Pervasive Agents
@@
Agents interpret user commands for device control via dynamic planning. For instance SAGE agent achieve 76% success rate on complex queries (Rivkin et al., 2024). Multi-agent systems like CASIT enhance IoT deployments by coordinating sensor data and detecting anomalies, outperforming single agents in data-rich environments (Zhong et al., 2024). The MuRAL dataset aids socially aware LLM development for smart environments with richly annotated sensor data (Chen et al., 2025a). Frameworks like LLMind integrate LLMs with AI modules for multi-device orchestration via natural language commands translated into device control scripts (Cui et al., 2024).
@@
Agents automate smartphone tasks by interpreting user instructions and executing page contents actions, eliminating manual interaction. AutoDroid (Wen et al., 2024a) use UI understanding and exploration for autonomous task completion. InfiGUIAgent (Liu et al., 2025) employs human-like reasoning for robust UI interaction. MobileGPT (Lee et al., 2024) prioritizes efficiency with adaptable task memory and hybrid failure recovery. Benchmarks like Android Agent Arena (A3) (Chai et al., 2025), LlamaTouch (Zhang et al., 2024g), and MobileAgentBench (Wang et al., 2024a) facilitate agent evaluation, while MobileSafetyBench (Lee et al., 2024) focuses on safe handling of sensitive operations.
@@
Agents provide analysis of medical and industrial data. AutoHealth (Cardenas et al., 2024) integrates agents into wearables for Parkinson’s monitoring. SpeziLLM (Zagar et al., 2025) enables privacy-preserving medical AI on local nodes. In Smart Manufacturing, IMVA (Lin et al., 2025) and MaViLa (Fan et al., 2025) offer multimodal reasoning for tasks like quality control and system orchestration, improving decision-making and automation.
@@
Agents enable intelligent vehicle dispatching (LiMeDa) (Chen et al., 2024b), realistic personal mobility generation (LLMob) (JIAWEI et al., 2024), and natural language interaction with smart building management for energy optimization and control (BuildingSage) (Dedeoglu et al., 2024). These systems demonstrate the transformative potential of LLM-based agents in creating intelligent and adaptive urban infrastructures.
@@ 5. Discussion
LLM-based agents represent a substantial advancement in Artificial Intelligence, enabling the tackling of highly complex tasks across diverse domains including software engineering, research, medicine, and robotics. As outlined in Section 2, these agents operate by perceiving their environment, reasoning and planning actions to achieve defined goals, and executing these plans through specific actions. This architecture typically uses a foundational model, either an LLM or a MLLM, and three essential modules: Memory, Planning, and Action. The Memory module stores necessary information, the Planning module designs action sequences, and the Action module generates individual steps. Pervasive computing stands out as a field for a practical application of this architecture. As detailed in Section 3, this domain focuses on enhancing human activity through integrated technology, providing computational power for a variety of applications that can significantly improve daily life. From smart homes and factories to the every day usage smartphones, embedding the computational power of agents within pervasive computing systems holds the potential to solve numerous challenges by autonomously handling previously intractable tasks. Section 4 introduced various strategies for integrating these agents into pervasive devices. These strategies are heavily influenced by the available resources on the deployment devices, and have facilitated the application of agents in pervasive computing, including personal assistants for health monitoring and device interaction, as well as building and traffic management in smart cities. However, several critical challenges remain under active research and warrant careful consideration.
Small/Quantized Models Limitations
A primary challenge is the performance degradation associated with the use of small or quantized models. As discussed in Section 4.1, pervasive computing scenarios often involve devices with limited computational power, memory, and energy, driven by connectivity constraints or the need for on-device data privacy. This frequently necessitates the deployment of agents relying on Small Language Models (SLMs) or quantized models, which can exhibit a significant performance drop compared to larger, more robust alternatives. To address this, researchers are exploring post-training alignment methods like Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) to better tailor models to specific application domains, as well as techniques for distributing model computation across multiple edge nodes. Furthermore, even when employing larger models, the inherent limitations in LLM architecture pose challenges. For instance the issue of hallucinations, that is the generation of highly probable but irrelevant or factually inconsistent token sequences, poses a significant obstacle to the proper functioning of agents. Ongoing research seeks to mitigate this through Retrieval-Augmented Generation (RAG) systems, which inject relevant external information into the model’s context, and post-training domain alignment techniques to better guide output coherence.
Generalization Problem
Another recognized limitation is the difficulty that LLMs face with generalization. As demonstrated in (Mirzadeh et al., 2024), even superficial changes like variable name alterations in mathematical problems can significantly degrade model performance, highlighting their reliance on training data patterns. This lack of robust generalization can be problematic for agents, particularly during the planning phase, potentially leading to suboptimal solutions or failure to follow task constraints, as emphasized in (Kambhampati et al., 2024). This study indicates that even advanced models like GPT-4o and Claude-3-opus struggle with planning tasks despite the use of techniques like ReAct (Yao et al., 2023) and Chain-of-Thought (Wei et al., 2022). To address this, modular LLM frameworks that incorporate critique and reformulation mechanisms, as proposed in (Gundawar et al., 2024), show promising results. Additionally, enhancing the reasoning capabilities of models at inference time has led to the development of Large Reasoning Models (LRMs). These models, during inference, first engage in a reasoning phase, similar to ”human-like thought”, before providing an answer. This process appears to increase the likelihood of the model to generate a correct response. Various methodologies are being explored in this area, including forcing the model to produce rationales (”thinking”) before giving the final answer (Zelikman et al., 2024), the use of Outcome-Supervised Reward Models (ORM) and Process Reward Models (PRM) to evaluate reasoning during training (Hosseini et al., 2024; Zhang et al., 2024c; Lightman et al., 2023), and fine-tuning techniques that enable models to detect and self-correct errors using rationales (Qu et al., 2024). The emergence of powerful LRMs like OpenAI’s o1, o3, o4-mini, Anthropic’s Claude 3.7 Sonnet, and DeepSeek’s V3 underscores the active research and development in this area.
@@
The external nature of the memory module in current agent architectures also presents a significant limitation. As agents interact with their environment, their experiences must be stored externally and then selectively reintroduced into the LLM’s context window. The limited size of this context window poses a constraint, particularly for agents requiring numerous interactions to achieve a goal, potentially leading to the loss of crucial information if earlier interactions are simplified or overly summarized, as discussed in Sections 2.2.2 and 4.3.1. To overcome this, researchers are exploring strategies to find better summarization methods, or to expand the context window, as seen with the 1 Million context window Qwen’s Qwen2.5-Turbo (Yang et al., 2025b) and Google’s Gemini 2.5 Pro. Another promising direction involves the direct integration of memory modules within the models themselves, such as the approach proposed by Google researchers in (Behrouz et al., 2024).
@@
The energy issue is particularly significant when we are facing with agents that have to use frequently LLMs for reasoning, planning, action, and memory management tasks. This challenge must be addressed especially in pervasive environments, where there is a wide heterogeneity of devices, from large cloud high resources computing servers to smartphones that have limited energy resources. As discussed in Section 4.1, an LLM has a very high energy cost per billion parameters, which leads to rapid battery drain on edge devices and high electricity consumption when used on fog or cloud servers. To mitigate this issue, recent works have explored multiple strategies to reduce energy consumption without severely compromising performance. In (Yin et al., 2024) the authors implemented a KV cache compression and swapping method to avoid the recomputation of the KV cache by the model. In (Maliakel et al., 2025) the researchers show that selecting models based on task complexity, and applying hardware-level techniques such as Dynamic Voltage and Frequency Scaling (DVFS), can cut energy usage by up to 50% without major accuracy loss. Another approach involves decentralized inference, where LLM layers are distributed across edge devices equipped with energy harvesting (e.g., solar panels). This allows sustained inference under energy constraints by dynamically scheduling computation based on device energy availability and predicted energy inflow (Khoshsirat et al., 2024). Furthermore, in (Samsi et al., 2023) the researchers tested multi-GPU inference and highlighted the substantial energy benefits of optimized LLM sharding strategies. In particular, using power capping strategies (limiting Watts usage), that increase the average inference time, save a lot energy without impacting the accuracy of the models. Finally, the GREEN-CODE (Ilager et al., 2025) framework introduces dynamic early exiting during inference using reinforcement learning. This enables LLMs to terminate inference early at intermediate layers when sufficient confidence is obtained, achieving energy reductions of $23$-$50$% in code generation tasks while maintaining output quality.
@@
Finally, is fundamental to consider privacy in pervasive computing due to the close interaction with users and the consequent handling of large amounts of sensitive data. Agents operating in this context must manage this information while guaranteeing the user’s complete privacy. For instance, when an agent interacts with applications on a smartphone or PC, it must exercise extreme caution in handling sensitive user data, including full names and the contents of private documents or files. The literature features various analyses and several benchmarks specifically designed to evaluate the security of agents in ensuring the preservation of user privacy. The PrivacyLens framework (Shao et al., 2024a) provides a multi-level evaluation of LLM agents’ awareness of contextual privacy norms. It introduces a pipeline that converts privacy-sensitive seeds into expressive vignettes and executable agent trajectories to uncover instances of private information leakage, such as sharing job-seeking details inappropriately, even with privacy-preserving prompts. PrivacyLens reveals that even advanced LLMs like GPT-4 leak sensitive information in a significant number of cases (up to 25.68%). AgentDojo (Debenedetti et al., 2024) assesses agents robustness against prompt injection attacks. This framework puts LLM agents with realistic scenarios like email management or banking app navigation while defending against malicious inputs aimed at extracting user data. AgentDojo highlights the persistent vulnerability of agents to adversarial inputs and underscores the need for privacy-aware defenses. Finally, Agent-SafetyBench (Zhang et al., 2024b) offers a broader evaluation of agents’ behavior across diverse environments, identifying privacy-related safety risks like unintentional data leakage. Evaluation across 2,000 test cases revealed that no agent achieved a safety score above 60%, highlighting a systemic lack of robustness and risk awareness in current LLM-based agent implementations.
Given these real privacy challenges that agents have to face, various strategies have been proposed to try to mitigate this issue. For instance, AgentDojo (Debenedetti et al., 2024) incorporates and evaluates several defense mechanisms, including secondary attack detection modules like Data delimiters, Prompt injection detection, Prompt sandwiching, and Tool Filter. These strategies significantly reduce the success rate of prompt injection attacks, demonstrating a drop to 7.5% when a tool filter is active. However, no single defense mechanism offers complete protection against these vulnerabilities. AutoDroid (Wen et al., 2024a) employs a Privacy Filter, a Personal Identifiable Information (PII) scanner, to detect sensitive information (such as names, phone numbers, and email addresses) within the prompt. This filter replaces any identified sensitive data, aiming to safeguard user privacy on smartphone devices. Furthermore, a layered defense strategy proposes three firewalls for agent and user data protection reducing private leakage from 70% to less than 2% (Abdelnabi et al., 2025): i) a data firewall isolates private data in environment outputs using task-relevant rules, without accessing the conversation; ii) a trajectory firewall checks the agent’s response against security rules post-decision, with an option to regenerate safer outputs; iii) an input firewall sanitizes external inputs by converting them to structured formats like JSON, removing manipulative language to reduce attack risks.

### v2

@@ 5.3.3. Action
  • Tool Use / API Calls: Agents can leverage external services or system functionalities by using specific tools or making API calls. This includes a wide range of interactions, such as operating smartphone applications, sending emails, querying databases, retrieving weather information, or controlling external devices. For example, in (wen2024autodroidv1), the agents possesses a defined set of possible actions (CLICK, CHECK/UNCHECK, SCROLL¡DIRECTION¿, INPUT¡TEXT¿) to interact with a smartphone device’s interface. In (rivkin2024aiot), the agent flexibly manages smart home devices using a dedicated device interaction tool. The API calls are executed using the SmartThings REST API for reading device attributes and sending commands, with an automatic error correction mechanism in case of call failures. Similarly, in (li2024ids) the agent is equipped with a series of tools, including Knowledge Retrieval, Data Extraction, Classification, and Long-Term Memory Retrieval, which it can use for intrusion detection analysis. These tools are invoked using a JSON format that specifies an action name (the tool’s identifier) and an action input (the associated settings or parameters for that tool).
@@
  • Code Execution: Agents may generate and execute code snippets to test conditions or directly interact with underlying systems. For example, in (wen2024autodroid), the agent interacts with the environment by translating a user’s natural language task into executable Python-like scripts. These scripts are generated by a small language model based on a structured app document and are then executed on the device to manipulate page contents elements, such as tapping buttons or scrolling content. Also, in (rao2024eco), a large language model is employed to generate Python code that guides the placement of microservices between edge and cloud infrastructure, based on real-time workload and latency data. This enables the LLM to adaptively recommend whether a microservice should run on the edge or in the cloud, with new placement decisions being automatically integrated into the running application.
@@
  • User Messaging (Natural Language Interaction): Agents can also act by directly interacting with users through natural language messages, providing support as intelligent assistants. For instance, in (fan2025mavila), the agent communicates its analysis of manufacturing images to users, performing anomaly detection and scene understanding to provide comprehensive support to smart factory employees. Also, in (zagar2025dynamic), the system uses a unified interface to interact with users naturally via chat, enabling secure, real-time processing of health data in healthcare domains such as Electronic Health Record (EHR) analysis, patient data explanation, and medical form automation.
5.4. Evaluations Metrics
The evaluation of LLM-based agents in pervasive computing environments must consider various variables such as the strict constraints of edge hardware or user personalization.
Resource consumption
Must be quantified to evaluate the cost of usage of an LLM Agent in resource contraint devices. Energy consumption (zhao2024edge) is a critical metric for battery-powered devices. It measures the Joules expended per inference task or per token generated. Studies on mobile agents often track the Battery Drain Rate (%/hour) during active agent operation compared to idle states. Also some measure track the inference latency such as Time Per Output Token (TPOT) for generative tasks (shao2025ai). For shared services, Context Switching Latency measures the overhead of loading different user profiles or agent personas into active memory on edge nodes (yin2024llm). Finally Memory Footprint is used to track the RAM usage of the agent, specifically the Peak Memory Usage during inference and the KV Cache Size (yin2024llm).
Contextual Accuracy and Personalization
Agents in pervasive environments must bridge the gap between ambiguous sensory inputs and specific user needs, requiring them to infer intent and adapt to personal preferences in real-time. In this case Implicit Preference Accuracy measures the agent’s ability to predict the correct item recommendations. Element Accuracy (Ele.Acc) measures the correctness of predicted UI coordinates on mobile screens, while Intersection over Union (IoU) evaluates the precision of bounding boxes when the agent must visually identify IoT devices or physical objects (zhang2024smartagent). Finally Intent Resolution (rivkin2024aiot) measures the agent’s success in disambiguating vague voice commands (e.g., ”make it cozy”) into specific device actions (e.g., ”dim lights to 50%, set temperature to 22°C”).
5.5. Applications of Pervasive Agents
@@
Agents interpret user commands for device control via dynamic planning. For instance SAGE agent achieve 76% success rate on complex queries (rivkin2024aiot). Multi-agent systems like CASIT enhance IoT deployments by coordinating sensor data and detecting anomalies, outperforming single agents in data-rich environments (zhong2024casit). The MuRAL dataset aids socially aware LLM development for smart environments with richly annotated sensor data (chen2025mural). Frameworks like LLMind integrate LLMs with AI modules for multi-device orchestration via natural language commands translated into device control scripts (cui2024llmind).
@@
Agents automate smartphone tasks by interpreting user instructions and executing page contents actions, eliminating manual interaction. AutoDroid (wen2024autodroidv1) use UI understanding and exploration for autonomous task completion. InfiGUIAgent (liu2025infiguiagent) employs human-like reasoning for robust UI interaction. MobileGPT (lee2024mobilegpt) prioritizes efficiency with adaptable task memory and hybrid failure recovery. Benchmarks like Android Agent Arena (A3) (chai2025a3), LlamaTouch (zhang2024llamatouch), and MobileAgentBench (wang2024mobileagentbench) facilitate agent evaluation, while MobileSafetyBench (lee2024mobilegpt) focuses on safe handling of sensitive operations.
@@
Agents provide analysis of medical and industrial data. AutoHealth (cardenas2024autohealth) integrates agents into wearables for Parkinson’s monitoring. SpeziLLM (zagar2025dynamic) enables privacy-preserving medical AI on local nodes. In Smart Manufacturing, IMVA (lin2025generative) and MaViLa (fan2025mavila) offer multimodal reasoning for tasks like quality control and system orchestration, improving decision-making and automation.
@@
Agents enable intelligent vehicle dispatching (LiMeDa) (chen2024llm), realistic personal mobility generation (LLMob) (jiawei2024large), and natural language interaction with smart building management for energy optimization and control (BuildingSage) (dedeoglu2024buildingsage). These systems demonstrate the transformative potential of LLM-based agents in creating intelligent and adaptive urban infrastructures.
@@ 6. Discussion
LLM-based agents represent a substantial advancement in Artificial Intelligence, enabling the tackling of highly complex tasks across diverse domains including software engineering, research, medicine, and robotics. As outlined in Section 2, these agents operate by perceiving their environment, reasoning and planning actions to achieve defined goals, and executing these plans through specific actions. This architecture typically uses a foundational model, either an LLM or a MLLM, and three essential modules: Memory, Planning, and Action. The Memory module stores necessary information, the Planning module designs action sequences, and the Action module generates and applies individual steps. Pervasive computing stands out as a field for a practical application of this architecture. As detailed in Section 3, this domain focuses on enhancing human activity through integrated technology, providing computational power for a variety of applications that can significantly improve daily life. From smart homes and factories to the every day usage smartphones, embedding the computational power of agents within pervasive computing systems holds the potential to solve numerous challenges by autonomously handling previously intractable tasks. Section 5 introduced various strategies for integrating these agents into pervasive devices. These strategies are heavily influenced by the available resources on the deployment devices, and have facilitated the application of agents in pervasive computing, including personal assistants for health monitoring and device interaction, as well as building and traffic management in smart cities. However, several critical challenges remain under active research and warrant careful consideration.
Small and Quantized Models Limitations
A primary challenge in pervasive computing is the resource intensity of LLMs relative to edge hardware constraints. As discussed in Section 5.1, deployment often necessitates the use of Small Language Models (SLMs) or aggressive quantization to meet energy and privacy requirements. However, a critical research gap specific to agentic AI is the degradation of reasoning action capabilities. When models have less parameters, their ability to maintain long-horizon goals, adhere to complex tool-use syntax, or self-correct during multi-step planning fails significantly more often than larger counterparts. Furthermore, smaller models are more prone to hallucinations, generating plausible but factually incorrect token sequences, which can lead to catastrophic failure in autonomous decision-making loops.
To address these limitations, future research must move beyond general knowledge distillation. The field requires specialized reasoning distillation, where SLMs are trained specifically to retain agentic behaviors (e.g., planning logic, API error handling). Additionally, research should focus on dynamic compute allocation frameworks that can assess task complexity in real-time, solving easy questions locally while dynamically offloading reasoning-heavy questions to larger remote models only when necessary. Finally, optimizing retrieval-augmented generation (RAG) specifically for low-parameter models is crucial to mitigate hallucinations without increasing model size.
@@
The external nature of the Memory module in current agent architectures also presents a significant limitation. As agents interact with their environment, their experiences must be stored externally and then selectively reintroduced into the LLM’s context window. The limited size of this context window poses a constraint, particularly for agents requiring numerous interactions to achieve a goal, potentially leading to the loss of crucial information if earlier interactions are simplified or overly summarized, as discussed in Sections LABEL:sec:memory and 5.3.1. To overcome this, researchers are exploring strategies to find better summarization methods, or to expand the context window, as seen with the 1 Million context window Qwen’s Qwen2.5-Turbo (yang2025qwen2) and Google’s Gemini 2.5 Pro. Another promising direction involves the direct integration of memory modules within the models themselves. For instance, in (behrouz2024titans), the Google researchers propose a novel architecture incorporating a neural long-term memory module. This module learns to memorize historical context at test time, demonstrating superior effectiveness over both Transformers and recent linear recurrent models, particularly in long context scenarios. Research should therefore concentrate on optimizing these on-device neural memory techniques and developing advanced context summarization algorithms that can maintain an effective agent execution trajectory without rapidly consuming the available context window.
@@
The Energetic issue is particularly significant when we are facing with agents that have to use frequently LLMs for reasoning, planning, action, and memory management tasks. This challenge must be addressed especially in pervasive environments, where there is a wide heterogeneity of devices, from large cloud high resources computing servers to smartphones that have limited energy resources. As discussed in Section 5.1, an LLM has a very high energy cost per billion parameters, which leads to rapid battery drain on edge devices and high electricity consumption when used on fog or cloud servers. The research challenge is how to achieve autonomous function while minimizing the energy cost of the continuous perceive-plan-act loop.
To mitigate this, future research must focus on optimizing both the core inference process and the agent’s decision cycle. Techniques that reduce computation for every inference step are vital: this includes optimizing internal structures through methods like KV cache compression and swapping to avoid expensive recomputations (yin2024llm), and adopting optimized LLM sharding strategies coupled with hardware techniques like power capping to improve energy per task (samsi2023words). Furthermore, research must target the agent’s decision-making process itself. This involves developing dynamic early-exiting inference frameworks like GREEN-CODE (ilager2025green), which use reinforcement learning to terminate inference when sufficient confidence is obtained, achieving substantial energy reductions. Another key direction is implementing event-triggered planning architectures that use low-power sensors to detect semantic events and only activate the high-power LLM reasoning engine when a proactive, high-value action is necessary. This prevents continuous, costly reasoning for trivial or irrelevant environmental changes.
@@
Finally, in pervasive computing, it’s fundamental to prioritise privacy because of its close interaction with users and the resulting handling of large amounts of sensitive data. Agents operating in this context must manage this information while guaranteeing the user’s complete privacy. For instance, when an agent interacts with applications on a smartphone or PC, it must exercise extreme caution in handling sensitive user data, including full names and the contents of private documents or files. The literature features various analyses and several benchmarks specifically designed to evaluate the security of agents in ensuring the preservation of user privacy. The PrivacyLens framework (shao2024privacylens) provides a multi-level evaluation of LLM agents’ awareness of contextual privacy norms. It introduces a pipeline that converts privacy-sensitive seeds into expressive vignettes and executable agent trajectories to uncover instances of private information leakage, such as sharing job-seeking details inappropriately, even with privacy-preserving prompts. PrivacyLens reveals that even advanced LLMs like GPT-4 leak sensitive information in a significant number of cases (up to 25.68%). AgentDojo (debenedetti2024agentdojo) assesses agents robustness against prompt injection attacks. This framework puts LLM agents with realistic scenarios like email management or banking app navigation while defending against malicious inputs aimed at extracting user data. AgentDojo highlights the persistent vulnerability of agents to adversarial inputs and underscores the need for privacy-aware defenses. Finally, Agent-SafetyBench (zhang2024agent) offers a broader evaluation of agents’ behavior across diverse environments, identifying privacy-related safety risks like unintentional data leakage. Evaluation across 2,000 test cases revealed that no agent achieved a safety score above 60%, highlighting a systemic lack of robustness and risk awareness in current LLM-based agent implementations.
While current mitigation strategies like Personal Identifiable Information (PII) scanner (e.g., AutoDroid (wen2024autodroidv1)) and layered firewalls (abdelnabi2025firewalls) have shown promise in reducing leakage rates, they often rely on static rules that can be avoided. Future research must evolve towards action-aware semantic firewalls, that do not merely filter text but analyse the intent and potential consequences of tool calls. For example, a firewall should dynamically restrict high-stakes API calls based on geolocation or user presence. Furthermore, research is needed into cryptographic verification methods that ensure critical physical actions originate from a verified human intent rather than an autonomous agent hallucination, creating a hard ”human-in-the-loop” requirement for high-risk interventions.